OpenStudy (anonymous):

Find the mean of the data: 14, 9, 2, 13, 16, 3, 13, 10. A. 10 B. 11.5 C. 13 D. 14.5

OpenStudy (anonymous):

I did this numerous times and my answer doesn't match any of thse

OpenStudy (anonymous):

these*

OpenStudy (zale101):

This is so easy, all u need to do is add up all the numbers, then divide by how many numbers there are. What did u get for ure answer?

OpenStudy (anonymous):

8.625

OpenStudy (anonymous):

actually its 10

OpenStudy (whpalmer4):

14+9+2+13+16+3+13+10 = 80, right?

OpenStudy (anonymous):

well that's what I was going to put since it the closest so I guess I had it right in a way

OpenStudy (anonymous):

yes its 80 and the mean is 10 @whpalmer4

OpenStudy (anonymous):

yep

OpenStudy (whpalmer4):

@agentenigma007 yes, I know the mean is 10, I'm trying to help @nicc96 figure out where they went astray

OpenStudy (whpalmer4):

80/8 = 10, which is the answer. it's pretty important to be able to do arithmetic correctly!

OpenStudy (anonymous):

I thought I was supposed to add all the numbers I had and divide them?

OpenStudy (whpalmer4):

Yes. 14,9,2,13,16,3,13,10 is all of the numbers, right? added up, they total 80. There are 8 of them, so we divide 80/8 to get the mean.

OpenStudy (anonymous):

80/8 cause the total of the 8 no.s which are there is 80

OpenStudy (anonymous):

i guess didn't add them correctly

OpenStudy (anonymous):

well thanks

OpenStudy (anonymous):

welcome :)

OpenStudy (whpalmer4):

sometimes when adding a string of numbers by hand, it is easier to reorder them, or add them in clumps. for example, in this string of numbers, the first number + the second two add up to 25. then I might add the 10 on the end to make 35. add the 3 to make 38, then add 16 to get 54, and 2*13 = 26 so 54+26 = 80. obviously, you need to cross numbers off the list when you use them! I find I make fewer mistakes if I avoid things like 8+3. Pay attention to where you make mistakes, so you can avoid them in the future!